OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (2, 0) and a directrix of y = −12?

OpenStudy (jdoe0001):

hmmm can you draw those 2 fellows, that is, the focus point and the directix?

OpenStudy (anonymous):

f(x) = negative one twenty fourth x squared plus one sixth x minus thirty five sixths f(x) = x2 − 4x − 35 f(x) = one twenty fourth x squared minus one sixth x minus thirty five sixths f(x) = −x2 + 4x + 35

OpenStudy (anonymous):

sorry i forgot to put the answers for you

OpenStudy (anonymous):